Translate the following mathematical statements into mathematical equations and solve unknown.
1. The volume of gas varies directly as the temperature and inversely as the pressure. If the volume is 250 cubic centimeters when the temperature is 330°K and the pressure is 25 pounds per square centimeter, what is the volume when the temperature is 230°K and the pressure is 20 pounds per square centimeter?

Answer by Theo(13342)

use combined variation formula.

formula becomes:

v = k * t / p

v is the volume
k is the constant of variation
t is the temperature
p is the pressure.

you are given that the volume is 250 when the temperature is 330 and the pressure is 25.

use that information to solve for k.

the formula becomes:

250 = k * 330 / 25

solve for k to get:

k = 250 * 25 / 330 = 18.93939393....

use that value of k to solve for v when t = 230 and p = 20.

you get:

v = 18.93939393..... * 230 / 20 = 217.8030303.....

that's your solution.